# Johannes Matzat        Aiko Schmeißer

### v4

@@ 7 Conclusion
Labor unions employ enormous resources to shape labor policies and welfare regulations through political activities such as lobbying legislators or supporting candidates financially. However, lasting change requires changes in the political preferences of the electorate. To understand the political power of labor unions, it is crucial to examine whether they can mobilize and influence the ideologies of millions of individuals at the unionized workplace. This paper analyzes the political effects of unionization, building on an establishment-level dataset that combines union elections with campaign contributions from employees spanning the 1980-2016 period in the United States. Comparing establishments that won and lost the union election in a stacked DiD model, we find that unionization increases contributions to Democratic candidates relative to Republican candidates by 12 percentage points for workers and 20 percentage points for managers, while we do not find a permanent impact on the overall amount of contributions. Overall, we show that labor unions influence the political preferences not only of union members but also of their firms’ management.
The results are indicative of a reduction of worker-manager cleavages in ideological positions, which is consistent with an improvement in workplace labor relations. If unionization fosters bargaining and communication between workers and managers on a more equal playing field, contact theory suggests an enhancement in managers’ understanding of workers’ political preferences. While the results may appear surprising in light of the strong opposition of employers toward unions in the United States, a distinction between ex-ante beliefs and ex-post effects of unionization seems crucial. The literature has found little evidence that unionization leads to higher wages (DiNardo and Lee, 2004; Frandsen, 2021; Freeman and Kleiner, 1990b) or reduced productivity (Dube et al., 2016; Sojourner et al., 2015), which could lower firms’ profitability. We welcome future work that studies more closely how managers form beliefs about unionization.
@@
Our findings may have implications for broader developments in U.S. politics. The longstanding decline in private-sector union density implies that millions of individuals have forfeited the engagement with unions, which has led to lasting shifts in political preferences. The erosion of unionization can be an important contribution to the increased alignment of workers with the political right that has been observed over the last decades (Gethin et al., 2022). More recently, prominent examples of strikes and union petition drives in Starbucks shops, Amazon warehouses, and healthcare facilities suggest a moment of resurgence for labor organization. Whether this trend persists may be consequential for the balance of political power and support for pro-labor politics in the United States.
@@ Appendix A Additional Figures and Tables
[…]
Figure A.3: Seasonally Adjusted Fatal Work Accidents, 1984-2012
@@
Figure A.5: Effect of Unionization on Total Contribution Amounts - DiD-RDD Results
@@
Figure A.6: Effect of Unionization on Democratic versus Republican Support - Vote Share Heterogeneity in Pre- versus Post-Effects
@@
Figure A.8: Individual-levels Effects - Event Study Results
@@
[A]: Extensive-margin value $x=0$
@@
[B]: Extensive-margin value $x=0.1$
@@
[C]: Extensive-margin value $x=0.7$
@@
[D]: Extensive-margin value $x=1$
@@
[E]: Extensive-margin value $x=3$
@@
Notes: The table presents results under different valuations of extensive and intensive margin effects. Reported are the DiD coefficients estimated in model (1) for the effect of unionization on the total amount contributed (columns (1) - (3)) and on the difference between the amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Amounts are transformed by the function that sets $m(\$)=\ln(\$)\,\text{if}\,\$>0$ and $m(0)=-x$. Panels A to E present results for varying the extensive-margin value $x$. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=33,103$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[C]: Amount1/4
$\delta_{\textrm{DiD}}$0.002 0.020 -0.032 0.219∗∗∗ 0.095∗∗∗ 0.178∗∗∗
@@
[D]: Untransformed amounts
$\delta_{\textrm{DiD}}$27.618 2.414 -22.951 116.721∗∗∗ 15.581∗∗ 65.375∗∗∗
@@
[E]: Borusyak, Jaravel, and Spiess (2021)
@@
[F]: Callaway and Sant’Anna (2021)
@@
[G]: Sample restriction: at least one matched contribution in pre-period
$\delta_{\textrm{DiD}}$0.008 0.050 -0.132 0.311∗∗ 0.223∗∗∗ 0.292∗∗∗
@@
[H]: Sample restriction: period between first and last matched contribution
@@
Notes: The table presents robustness checks for our DiD estimates of the effect of unionization on the total amount contributed (columns (1) - (3)) and on the difference between the amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panel A shows the baseline results from the stacked DiD model (1) with IHS-transformed amounts. In Panels B and C, outcomes are transformed as log(amount+1) and amount1/4, respectively, while in Panel D we use untransformed amounts after winsorizing at the top and bottom percentile. Panel E presents results from the imputation approach introduced by Borusyak et al. (forthcoming), and Panel F implements the DiD estimator of Callaway and Sant’Anna (2021), where we use both never-treated establishments (i.e., lost elections) and not-yet-treated establishments (i.e., won elections in later cycles) as comparison units. In Panels A to F, $N=33,103$ establishment-cycle observations. In Panel G, we constrain the sample to establishments for which we have matched at least one contribution in the three cycles before the union election ($N=15,792$). Panel H restricts the observation window for each establishment to the cycles between the first matched contribution before the union election and the last matched contribution after the union election ($N=17,911$). Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[C]: Supervisor tasks “very important” (4 out of 5 in ranking)
@@
[D]: Non-managerial supervisors as workers
@@
[…]
@@
[C]: Local candidates
$\delta_{\textrm{DiD}}$0.047 0.024 -0.034 0.158∗∗∗ 0.045 0.130∗∗∗
@@
Table A.9: Effects of Losing a Union Election
IHS(\$ to all candidates) IHS(\$ to Dem.) -- IHS(\$ to Rep).
@@
[A]: Stacking
$\delta_{\textrm{DiD}}$0.049 -0.026 0.070 0.057 -0.013 0.037
(0.088) (0.040) (0.053) (0.097) (0.043) (0.057)
N 31,501 31,501 31,501 31,501 31,501 31,501
[B]: Borusyak, Jaravel, and Spiess (2021)
$\delta_{\textrm{DiD}}$0.048 -0.029 0.075 0.080 -0.007 0.048
(0.090) (0.045) (0.059) (0.100) (0.049) (0.064)
N 16,658 16,658 16,658 16,658 16,658 16,658
[C]: Callaway and Sant’Anna (2021)
$\delta_{\textrm{DiD}}$0.043 -0.038 0.061 0.076 -0.007 0.053
(0.095) (0.047) (0.064) (0.105) (0.052) (0.070)
N 16,658 16,658 16,658 16,658 16,658 16,658
Notes: The table presents DiD estimates for the effect of losing a union election versus holding no election. We compare establishments with a lost union election in a given cycle (treated cohort) with establishments with a lost union election in one of the next two cycles (control cohorts) in a DiD design. Thereby, we estimate short-term effects of losing an election (i.e., for event times $k=\{0,1\}$). Panel A shows results from a stacked DiD model, and Panels B and C implement the staggered DiD estimators of Borusyak et al. (forthcoming) and Callaway and Sant’Anna (2021). See Appendix C for details of the implementation. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
IHS(\$ to all candidates) IHS(\$ to Dem.) -- IHS(\$ to Rep.)
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A]: At least one matched PAC donation one year before union election (baseline)
@@
[B]: At least one matched PAC donation two years before union election
$\delta_{\textrm{DiD}}$0.041 0.461 0.035 0.437∗∗∗ 0.652∗∗∗ 0.320
@@
[C]: At least one matched PAC or candidate donation one year before union election
@@
[D]: At least one matched PAC or candidate donation two years before union election

### v5

@@ 7 Conclusion
Labor unions employ enormous resources to shape labor policies and welfare regulations through political activities such as lobbying legislators or supporting candidates financially. However, lasting change requires changes in the political preferences of the electorate. To understand the political power of labor unions, it is crucial to examine whether they can mobilize and influence the ideologies of millions of individuals at the unionized workplace. This paper analyzes the political effects of unionization, building on an establishment-level dataset that combines union elections with campaign contributions from employees spanning the 1980-2016 period in the United States. Comparing establishments that won and lost the union election in a stacked DiD model, we find that unionization increases contributions to Democratic candidates relative to Republican candidates by 12 percentage points for workers and 20 percentage points for managers, while we do not find a lasting impact on the overall amount of contributions. Overall, we show that labor unions influence the political preferences not only of union members but also of their firms’ management.
We also document that unionization leads to larger ideological shifts in settings where labor-management relations are more cooperative, such as in union elections without unfair labor practice charges and where the employer and union reach a collective bargaining agreement. These patterns suggest an important role of workplace contact, in particular for explaining why managers align their political views with those of workers and the union. The contact hypothesis posits that cooperative interactions can enhance perspective-taking and reduce biases in beliefs about opposing groups that may prevail among managers. Biased beliefs may also help explain why many employers in the U.S. remain strongly opposed to unionization. The literature finds little evidence that unionization strongly increases wages (DiNardo2004; Frandsen2020; Freeman1990) or reduces productivity (Dube2016; Sojourner2015a), which could harm firms profitability; yet strong employer resistance persists. We welcome future research that explicitly studies how managers form beliefs about the impact of unionization and what factors sustain anti-union sentiment among employers.
@@
Our findings may have implications for broader developments in U.S. politics. The longstanding decline in private-sector union density implies that millions of individuals have forfeited the engagement with unions, which has led to lasting shifts in political preferences. The erosion of unionization can be an important contribution to the increased alignment of workers with the political right that has been observed over the last decades (Gethin2022). More recently, prominent examples of strikes and union petition drives in Starbucks shops, Amazon warehouses, and healthcare facilities suggest a moment of resurgence for labor organization. Whether this trend persists may be consequential for the balance of political power and support for pro-labor politics in the United States.
Figures and Tables
Figure 1: Trends in Contributions for Won and Lost Union Elections
Notes: The figure displays trends in average contribution amounts (in 2010 USD) from all employees in an establishment, separately by union election outcome and election cycle relative to the union election. The left graph shows the mean of IHS-transformed total contributions to all candidates. The right graph show the mean difference between IHS-transformed amounts contributed to Democratic and Republican candidates. $N=42,441$ establishment-cycle observations.
Figure 2: Effect of Unionization on Candidate Contributions
Notes: The figures report the event-study coefficients $\delta_{s}$ estimated in model (2). The sample includes all establishments with a pro-union vote share between 20% and 80% and covers three election cycles (six years) before and after the union election. $N=33,103$ establishment-cycle observations. Below each graph the DiD coefficient from model (1) is reported. In the graphs on the left side, the outcome is the IHS-transformed total amount contributed to all candidates. In the graphs on the right side, the outcome is the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates. Results are reported for contributions from all employees (top part), from non-managerial workers (middle part), and from managers and supervisors (lower part). 95% confidence intervals are depicted for standard errors clustered at the establishment level.
Figure 3: Vote Share Tests
(A) Vote Share Heterogeneity
(B) Vote Share Bandwidth Sample Restrictions
Notes: The graphs show RDD-motivated placebo and robustness tests for the effect of unionization on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates. Panel (a) reports the $\delta_{g}$ coefficients estimated in model (3). The vote share distribution is partitioned into six bins, indicated on the x-axis. The omitted reference group is 20-35%. Panel (b) reports DiD coefficients estimated in model (1). Each dot refers to a single DiD coefficient that is estimated among elections with a union vote share in a given bandwidth around the 50% cutoff. Estimates from smaller bandwidths compare changes between increasingly close elections. Results are always shown separately for contributions from non-managerial workers (“workers”) and from managers and supervisors (“managers”). 95% confidence intervals are depicted for standard errors clustered at the establishment level.
Table 1: Election and Contribution Descriptive Statistics
All Union Loss Union Win
[A] Election characteristics
Number of elections 6,063 3,397 2,666
Union vote share (average) .4950 .3204 .7175
Number of votes (average) 119.37 135.31 99.06
Number of votes (total) 723,752 459,661 264,091
[B] Contribution characteristics
Amount (total, in million 2010 USD) 105.82 65.38 40.43
Number of contributions (total) 357,436 204,797 152,639
Number of donors (total) 46,719 26,661 20,243
Number of recipients (total) 9,942 7,208 5,681
Notes: Data from NLRB union certification elections, which have at least one employee contribution matched in any of seven election cycles around the union election (three before, cycle of union election, three after). Contribution characteristics refer to the total numbers over all these seven election cycles.
Table 2: Contributions by Donor and Recipient
Recipient: Donor: All employees Workers Managers
All 2,493.24 313.80 1,339.38
Candidates 1,181.96 173.42 594.44
Democratic candidates 575.85 112.79 261.76
Republican candidates 586.98 56.61 320.66
Political action committees 1,311.28 140.38 744.94
Party PACs 364.92 52.52 192.77
Interest group PACs 937.22 86.37 549.31
Notes: The table reports mean values for the amount contributed in each of the 42,441 establishment-cycle combinations in the estimation sample. All amounts are in 2010 USD. Values are reported separately for contributions from all employees, from only non-managerial workers (“workers”), and from only managers and supervisors (“managers”). The difference in the amounts from all employees and the total from workers and managers is driven by contributions for which we were unable to classify the occupation.
Table 3: Composition versus Individual-Level Effects
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep.)
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A] Composition effects
$\delta_{\textrm{DiD}}$ -0.027 0.046 -0.067 0.071 0.053 0.037
(0.070) (0.036) (0.051) (0.064) (0.029) (0.045)
N 33,103 33,103 33,103 33,103 33,103 33,103
[B] Individual-level effects
$\delta_{\textrm{DiD}}$ 0.047 0.507 0.183 0.633∗∗∗ 0.727∗∗ 0.594∗∗
(0.162) (0.261) (0.260) (0.160) (0.275) (0.274)
N 22,799 5,243 13,104 22,799 5,243 13,104
Notes: The table reports DiD coefficients for the composition and individual-level effects of unionization on the IHS-transformed total amount contributed (columns (1) - (3)) and on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panel A shows results from the establishment-level model (1). The establishment outcomes for the post-election periods are computed from the pre-election contributions of donors matched to an establishment in the respective post-election period. Outcomes for the pre-election periods are constructed as before from the actual contributions in those periods. Panel B shows results from the individual-level model (6). The sample includes incumbent employees who made at least one contribution to a PAC in the year before the union election that is matched to the establishment. The individual $\times$ cycle panel aggregates all the contributions of these donors three cycles before to three cycles after the union election, irrespective of whether or not they are matched to the initial establishment. In the regressions, each incumbent employee is weighted by the inverse of the total number of donating incumbent employees in the establishment. All samples include establishments with a pro-union vote share between 20 and 80%. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 4: Differentiating Candidates by Within-Party Ideology
IHS(\$ to Democrats) IHS(\$ to Republicans)
Liberal Moderate Moderate Conservative
(1) (2) (3) (4)
[A] All employees
$\delta_{\textrm{DiD}}$ 0.119∗∗∗ -0.017 -0.0690 -0.153∗∗∗
(0.046) (0.054) (0.055) (0.049)
[B] Workers
$\delta_{\textrm{DiD}}$ 0.053 0.031 -0.015 -0.031
(0.030) (0.024) (0.022) (0.026)
[C] Managers
$\delta_{\textrm{DiD}}$ 0.088∗∗ 0.014 -0.056 -0.123∗∗∗
(0.035) (0.039) (0.040) (0.037)
Notes: The table reports DiD coefficients, estimated in model (1), for the effect of unionization on IHS-transformed amounts contributed to different candidate groups. Liberal (moderate) Democrats refer to Democratic candidates with a CF score below (above) the median CF score of all Democratic candidates observed in our sample of matched contributions. Conservative and moderate Republicans are differentiated accordingly using the median Republican CF score. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=33,103$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 5: Differentiating Candidates by Union Support
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep.)
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A] Candidates supported by union PAC
$\delta_{\textrm{DiD}}$ 0.065 0.030 0.047 0.100 0.047 0.102∗∗
(0.057) (0.030) (0.043) (0.057) (0.028) (0.042)
[B] Candidates not supported by union PAC
$\delta_{\textrm{DiD}}$ -0.052 0.008 -0.046 0.225∗∗∗ 0.075∗∗ 0.134∗∗∗
(0.070) (0.036) (0.053) (0.072) (0.033) (0.052)
Notes: The table presents DiD coefficients, estimated in model (1), for the effect of unionization on the IHS-transformed total amount contributed (columns (1) - (3)) and on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panels A and B distinguish contribution amounts to candidates supported versus not supported by PACs associated with each election’s union organization. See Appendix Table A.11 for a list of all union organizations and the party composition of their matched donations. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=29,757$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 6: Contributions to Political Action Committees
Party PACs Interest group PACs
All Dem - Rep All Union Member Corporation Trade Dem - Rep
orga. assoc.
(1) (2) (3) (4) (5) (6) (7) (8)
[A] All employees
$\delta_{\textrm{DiD}}$ -0.025 0.097∗∗ -0.082 0.022 -0.011 -0.093∗∗ -0.026 0.060
(0.052) (0.048) (0.064) (0.011) (0.031) (0.041) (0.044) (0.041)
[B] Workers
$\delta_{\textrm{DiD}}$ 0.062 0.010 0.088∗∗ 0.024∗∗∗ 0.044∗∗ -0.020 0.021 0.024
(0.032) (0.028) (0.035) (0.007) (0.019) (0.021) (0.016) (0.027)
[C] Managers
$\delta_{\textrm{DiD}}$ -0.001 0.102∗∗∗ -0.093 0.006 -0.002 -0.082∗∗ -0.026 0.081∗∗
(0.034) (0.031) (0.049) (0.007) (0.018) (0.034) (0.033) (0.032)
Notes: The table reports DiD coefficients, estimated in model (1), for the effect of unionization on IHS-transformed amounts contributed to different committee groups. In columns (2) and (8) the dependent variable is the difference between the IHS-transformed amounts contributed to Democratic and Republican committees. Interest group PACs are categorized as “Democratic” (“Republican”) if more (less) than 50% of their own campaign contributions goes to Democratic candidates. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=33,103$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 7: Heterogeneous Effects by Labor Relations Environment
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep.)
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A.1] States without Right-to-Work law
$\delta_{\textrm{DiD}}$ 0.045 0.066 -0.039 0.284∗∗∗ 0.131∗∗∗ 0.218∗∗∗
(0.090) (0.050) (0.067) (0.088) (0.046) (0.064)
N 26,208 26,208 26,208 26,208 26,208 26,208
[A.2] States with Right-to-Work law
$\delta_{\textrm{DiD}}$ -0.055 -0.119 0.008 0.016 0.070 0.142
(0.170) (0.082) (0.125) (0.177) (0.077) (0.117)
N 6,895 6,895 6,895 6,895 6,895 6,895
[B.1] Elections without unfair labor practice charge
$\delta_{\textrm{DiD}}$ 0.056 0.093 -0.052 0.347∗∗∗ 0.153∗∗∗ 0.290∗∗∗
(0.111) (0.062) (0.084) (0.110) (0.059) (0.079)
N 17,402 17,402 17,402 17,402 17,402 17,402
[B.2] Elections with unfair labor practice charge
$\delta_{\textrm{DiD}}$ 0.054 -0.044 0.001 0.167 0.102 0.132
(0.120) (0.065) (0.091) (0.121) (0.057) (0.085)
N 14,658 14,658 14,658 14,658 14,658 14,658
[C.1] Won elections with collective bargaining contract
$\delta_{\textrm{DiD}}$ 0.070 0.112 0.080 0.468∗∗∗ 0.207∗∗∗ 0.356∗∗∗
(0.104) (0.058) (0.077) (0.105) (0.053) (0.076)
N 26,089 26,089 26,089 26,089 26,089 26,089
[C.2] Won elections without collective bargaining contract
$\delta_{\textrm{DiD}}$ 0.002 -0.047 -0.109 0.037 0.054 0.076
(0.098) (0.053) (0.074) (0.095) (0.049) (0.067)
N 26,859 26,859 26,859 26,859 26,859 26,859
Notes: The table presents DiD coefficients, estimated in model (1), for the effect of unionization on the IHS-transformed total amount contributed (columns (1) - (3)) and on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panels A.1 and A.2 distinguish between establishments in states with versus without Right-to-Work laws in the union election year. Panels B.1 and B.2 differentiate union elections by whether the union filed an unfair labor practice charge (for violation of sections 8(a)(1), 8(a)(3), or 8(a)(4) of the NLRA) in the six months before or after the election. In Panels C.1 and C.2, we differentiate won union elections that have a contract expiration notice reported to the FCMS in the 5 years following the election and those that do not. In both C.1 and C.2, the control group consists of all lost elections. The sample includes establishments with a pro-union vote share between 20 and 80%. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Supplemental Appendix:
Do Unions Shape Political Ideologies at Work?
@@ Appendix A Additional Figures and Tables
Notes: The figure plots the density of union vote shares for all 6,063 union elections included in our matched estimation sample. The Frandsen2017 test strongly rejects continuity in the union vote share density at the 50% cutoff (p-value =.002=.002 for $k=0$ and p-value =.003=.003 for $k=.02$).
@@
Figure A.4: Vote Share Tests for Effect of Unionization on Total Contribution Amounts
@@
Figure A.5: Vote Share Heterogeneity in Pre- versus Post-Effects
@@
Figure A.7: Individual-level Effects - Event Study Results
@@
Table A.4: DiD Results
IHS(\$ donated to …)
All candidates Dem. Rep. Dem. - Rep.
(1) (2) (3) (4)
[A] All employees
$\delta_{\textrm{DiD}}$ 0.033 0.092 -0.147∗∗ 0.239∗∗∗
(0.079) (0.063) (0.065) (0.079)
[B] Workers
$\delta_{\textrm{DiD}}$ 0.026 0.073∗∗ -0.050 0.123∗∗∗
(0.043) (0.035) (0.032) (0.040)
[C] Managers
$\delta_{\textrm{DiD}}$ -0.019 0.074 -0.130∗∗∗ 0.204∗∗∗
(0.060) (0.047) (0.049) (0.056)
Notes: The table reports DiD coefficients, estimated in model (1), for the effect of unionization on employees’ campaign contributions. The outcome in column (1) is the IHS-transformed total amount contributed to all candidates. Columns (2) and (3) separately report IHS-transformed contributions to Democratic and Republican candidates, respectively, while column (4) reports the difference in IHS-transformed contributions to Democratic versus Republican candidates. Panel A includes contributions from all employees, Panel B focuses on non-managerial workers, and Panel C on managers and supervisors. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=33,103$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[A] Extensive-margin value $x=0$
@@
[B] Extensive-margin value $x=0.1$
@@
[C] Extensive-margin value $x=0.7$
@@
[D] Extensive-margin value $x=1$
@@
[E] Extensive-margin value $x=3$
@@
Notes: The table presents results under different valuations of extensive and intensive margin effects. Reported are the DiD coefficients estimated in model (1) for the effect of unionization on the total amount contributed (columns (1) - (3)) and on the difference between the amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Amounts are transformed by the function that sets $m(\mathdollar)=\ln(\mathdollar)\,\text{if}\,\mathdollar>0$ and m(0)=xm(0)=-x. Panels A to E present results for varying the extensive-margin value $x$. The sample includes establishments with a pro-union vote share between 20 and 80%. $N=33,103$ establishment-cycle observations. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[C] Amount1/4
$\delta_{\textrm{DiD}}$ -0.002 0.020 -0.032 0.219∗∗∗ 0.095∗∗∗ 0.178∗∗∗
@@
[D] Untransformed amounts
$\delta_{\textrm{DiD}}$ -27.618 2.414 -22.951 116.721∗∗∗ 15.581∗∗ 65.375∗∗∗
@@
[E] Borusyak, Jaravel, and Spiess (2021)
@@
[F] Callaway and Sant’Anna (2021)
@@
[G] Sample restriction: at least one matched contribution in pre-period
$\delta_{\textrm{DiD}}$ -0.008 0.050 -0.132 0.311∗∗ 0.223∗∗∗ 0.292∗∗∗
@@
[H] Sample restriction: period between first and last matched contribution
@@
Notes: The table presents robustness checks for our DiD estimates of the effect of unionization on the total amount contributed (columns (1) - (3)) and on the difference between the amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panel A shows the baseline results from the stacked DiD model (1) with IHS-transformed amounts. In Panels B and C, outcomes are transformed as log(amount+1) and amount1/4, respectively, while in Panel D we use untransformed amounts after winsorizing at the top and bottom percentile. Panel E presents results from the imputation approach introduced by Borusyak2021, and Panel F implements the DiD estimator of Callaway2021, where we use both never-treated establishments (i.e., lost elections) and not-yet-treated establishments (i.e., won elections in later cycles) as comparison units. In Panels A to F, $N=33,103$ establishment-cycle observations. In Panel G, we constrain the sample to establishments for which we have matched at least one contribution in the three cycles before the union election ($N=15,792$). Panel H restricts the observation window for each establishment to the cycles between the first matched contribution before the union election and the last matched contribution after the union election ($N=17,911$). Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[C] Supervisor tasks “very important” (4 out of 5 in ranking)
@@
[D] Non-managerial supervisors as workers
@@
[E] 80th percentile of census-tract median income
$\delta_{\textrm{DiD}}$ 0.059 -0.055 0.156∗∗ 0.216∗∗∗
(0.066) (0.065) (0.064) (0.062)
[F] 90th percentile of census-tract median income
$\delta_{\textrm{DiD}}$ 0.031 -0.018 0.243∗∗∗ 0.100
(0.071) (0.057) (0.070) (0.054)
Notes: The table presents robustness checks for alternative worker-manager classifications. Reported are the DiD coefficients estimated in model (1) for the effect of unionization on the IHS-transformed total amount contributed (columns (1) and (2)) and on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates (columns (3) and (4)). $N=33,103$ establishment-cycle observations. Panel A shows the baseline results in which “managers” are defined as donors in “Management occupations” (SOC group 11) or in occupations above the 80th percentile of supervisor tasks and independent judgment. “Workers” are all remaining donors with a classified occupation. In Panel B, we increase the cutoff for supervisor tasks and independent judgment to the 90th percentile. Panel C, instead, uses an absolute cutoff for the importance of supervisor tasks and independent judgment (both need to be “very important”, i.e., have a score of 4 or above in the 5-score ranking). In Panel D, we only consider “Management occupations” (SOC group 11) as “managers” and treat all other classified occupations as “workers” (including those with high importance in supervisor tasks and independent judgment). In Panels E and F, we define “managers” (“workers”) as individuals residing in a census tract with a median income above (below) the 80th or 90th percentile of the state-specific distribution of census-tract median incomes. See Appendix B.3 for more details on the classifications. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
[C] Local candidates
$\delta_{\textrm{DiD}}$ -0.047 0.024 -0.034 0.158∗∗∗ 0.045 0.130∗∗∗
@@
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep.)
@@
[A] At least one matched PAC donation one year before union election (baseline)
@@
[B] At least one matched PAC donation two years before union election
$\delta_{\textrm{DiD}}$ -0.041 0.461 0.035 0.437∗∗∗ 0.652∗∗∗ 0.320
@@
[C] At least one matched PAC or candidate donation one year before union election
@@
[D] At least one matched PAC or candidate donation two years before union election